A national standard requires that public bridges over 20 feet in length must be inspected and rated every 2 years. The rating scale ranges from 0​ (poorest rating) to 9​ (highest rating). A group of engineers used a probabilistic model to forecast the inspection ratings of all major bridges in a city. For the year​ 2020, the engineers forecast that 9​% of all major bridges in that city will have ratings of 4 or below. Complete parts a and b.
a. Use the forecast to find the probability that in a random sample of 7 major bridges in the​ city, at least 3 will have an inspection rating of 4 or below in 2020.

Answer by Boreal(15235)

Look at the probability 0, 1, 2 will have such an inspection rating, then subtract that from 1 to give the probability of having at least 3
for 0, it is 0.91^7=0.517
for 1 it is 7*0.91^6*0.09=0.358
for 2 it is 7C2*0.91^5*0.09^2=9=0.106
These three add to 0.981, so 1-0.981=0.019 ANSWER